suppose the ball was 2.74 feet above ground when Josh hit it, and that it reached a maximum height of approximately 116.3 feet when it was approximately 215.3 feet away from where he hit the ball. The ball lands after traveling a ground distance of approximately 433.1 feet. find the other root?
Answer by stanbon(75887)

suppose the ball was 2.74 feet above ground when Josh hit it, and that it reached a maximum height of approximately 116.3 feet when it was approximately 215.3 feet away from where he hit the ball. The ball lands after traveling a ground distance of approximately 433.1 feet. find the other root?
--------
Vertex: (215.3,116.3)
---
The x-coordinate of the vertex is at the midpoint between the roots.
One root is 433.1-215.3 = 217.8 to the right of 215.3.
The other root is 217.8 to the left of 215.3 or 215.3-217.8 = -2.5
---
left root 215.3-217.8 = -2.5
right root 215.3+217.8 = 433.1
================
Cheers,
Stan H.
===========